# No Dimensional Sampling Coresets for Classification

### v1

@@ 4.2 Coreset for Logistic Function
Proof  Simplifying notation, let $\phi(t)=\operatorname{{\rm logistic}}(t)$. For a given $\alpha>0$, if we set $t=\alpha z$ and $K=\alpha^{2}k$,
@@
Up to this point, we have established $\sup_{t\in[0,\ \infty)}f_{K}(t)\leq 25+\frac{75K}{4\log K}$ under the condition that $K\geq e$. To conclude the argument, we need to determine an upper bound for $\sup_{t\in[0,\ \infty)}f_{K}(t)$ when $0<K\leq e$. Utilizing Lemma 19, for $0<K\leq e$, we deduce
@@
Lemma 27
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. Then $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}_{\operatorname{{\rm logistic}}})$ and $(\mathcal{X},P,\bar{\mathcal{L}}_{\operatorname{{\rm logistic}}})$.
Proof  As $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, we only need to determine $s(y)$ for $y$ with $\|y\|_{H}\leq A$. Henceforth, we assume that $\|y\|_{H}\leq A$. For $w\in\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{A}\sqrt{\max(1,\log(A^{2}k))}\}$, Lemma 26 indicates that Lemma 20 is applicable with $B_{1}=A,B_{2}=\frac{1}{A}\sqrt{\max(1,\log(A^{2}k))},\phi(-B_{1}B_{2})\leq 1+\sqrt{\max(1,\log(A^{2}k))}=M$, and
@@
Using Lemma 21 with $\gamma(\|x\|_{H})=\frac{\operatorname{{\rm logistic}}(0)}{M\beta_{\operatorname{{\rm logistic}}}(\|x\|_{H})}\geq\frac{1}{2M\beta_{\operatorname{{\rm logistic}}}(\|x\|_{H})}$, we conclude that
@@
So, $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$. The proof for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$ proceeds in a similar manner.
Theorem 28 (No Dimensional Logistic Coreset)
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, $S=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$, and $C=(2A+1)\max(4Ak,2.5)+8Ak+1$. For $m\geq\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}\in\mathcal{X}$ according to $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\ell^{H}_{\operatorname{{\rm logistic}},k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm logistic}},k})$ ) with probability at least $1-\delta$.
Proof  Lemma 27 asserts that $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ serves as an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$ and $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm logistic}},k})$. Since it is a constant function, $s$-sensitivity sampling is equivalent to sampling according to $P$. It is worth noting that $\operatorname{{\rm logistic}}$ is a $1$-Lipschitz, convex, and decreasing function. For $E_{1}=A$, we have
@@
This implies that Definition 8 is satisfied for $\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k}$ and thus Theorem 10 concludes the statement for $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta}\right)$.
@@
Proof  Because of similarity, we only handle the case $(\mathbb{R}^{d},P,\mathbb{R}^{d},\ell_{\operatorname{{\rm logistic}},k})$. Given that Lemma 27 asserts the constant function $s(x)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ as an upper sensitivity function for $\mathcal{L}_{\operatorname{{\rm logistic}},k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$. To clarify, every function in $\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}}$ is a function in $\mathcal{L}_{\operatorname{{\rm logistic}},k}$ scaled by a positive constant. This implies ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ))={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)).$ We would like to remind that
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$ only includes half-spaces, empty set, and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  Matousek 2002). Leveraging Theorem 3, we derive the statement for $m=O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, thereby completing the proof.
In the previous theorem, we assumed that the data is hard-bounded. In the following results, we try to relax this assumption to a kind of soft bounding similar to what we have in Theorem 24.
Lemma 30
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}^{2}.$ Then
@@
Proof  As $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}^{2}\geq 2E^{2}_{1})\leq\frac{1}{2}$. For an arbitrary $y\in\mathcal{X}$, define $\mathcal{X}_{y}=\left\{x\in\mathcal{X}\colon\|x\|_{H}\leq\sqrt{2}E_{1}\max(1,\frac{\|y\|_{H}}{E_{1}})\right\}$. Additionally, define
@@
Lemma 26 indicates that Lemma 20 is applicable with
@@
(Lemma 20) $\leq\sup_{\|w\|_{H}\leq B_{2}}\frac{1}{\int_{x\in\mathcal{X}_{y}}\frac{\operatorname{{\rm logistic}}(0)}{M_{y}\beta(\|x\|_{H})}{\rm d}P(x)}$
@@
(*(*$\quad$( * ) $\leq 2\left(2+\frac{\|y\|_{H}}{E_{1}}\right)\sqrt{\max(1,\log(2E_{1}^{2}k)}\frac{340(1+kE_{1}^{2})}{\max(1,\log(2kE_{1}^{2}))}.$
@@
To see (***), note that $M_{y}\leq\left(2+\frac{\|y\|_{H}}{E_{1}}\right)\sqrt{\max(1,\log(2E_{1}^{2}k)}$, $\operatorname{{\rm logistic}}(2)\geq 0.5$, and
@@
$\sup_{\|w\|_{H}\geq B_{2}}\frac{\ell^{H}_{{}_{\operatorname{{\rm logistic}},k}}(y,w)}{\int_{x\in\mathcal{X}}\ell^{H}_{{}_{\operatorname{{\rm logistic}},k}}(x,w){\rm d}P(x)}$ $\leq\sup_{\|w\|_{H}\geq B_{2}}\frac{\frac{1}{k}\|w\|_{H}^{2}+\log(1+e^{-\boldsymbol{K}(y,w)})}{\frac{1}{k}\|w\|_{H}^{2}}$
$\leq\sup_{\|w\|_{H}\geq B_{2}}\left[1+\frac{k}{\|w\|_{H}^{2}}\log(1+e^{-\boldsymbol{K}(y,w)})\right]$
@@
A similar argument concludes the statement for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$.
Theorem 31
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\boldsymbol{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}$,
@@
Proof  Lemma 30 indicates that $s(\cdot)$ and $S$ are an upper sensitivity function its corresponding total sensitivity for both $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$ and $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$. Note that $\operatorname{{\rm logistic}}(\cdot)$ is decreasing and convex. using Equation (6), we have
@@
Now, the proof follows by applying Theorem 18.
Returning to the scenario where our Hilbert space is $\mathbb{R}^{d}$ equipped with the standard inner product as its kernel, to utilize Theorem 3, the computation of ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{{\mathcal{L}_{\operatorname{{\rm logistic}},k}}},\succ))$ is necessary. The following theorem serves as a valuable instrument for handling the VC-dimension of linked-range spaces.
@@
Theorem 32 (Anthony and Bartlett 2009, Theorem 8.14)
@@
The inequality $f_{w}(x)>r$ is not simply computable since it contains the evaluation $\|x\|$ (to compute $s(x)$), which cannot be done by the allowed operations. To circumvent this issue, we propose a modification to the upper sensitivity function $s$ outlined in Lemma 30 as follows:
@@
So, using Theorems 3 and 32, we have the next statement.
Theorem 33
@@
One might question the necessity of having Theorems 28 and 25 while Theorems 31 and 33 appear stronger and broader. It is crucial to note that in Theorems 31 and 33, the process demands re-sampling using $s$-sensitivity samples, which relies on evaluating the function $s$. In contrast, Theorems 28 and 25 permit the direct (via uniform sampling) use of given data points since they were originally sampled from $P$.
4.3 Coreset for svm Function
We would like to remind the reader that $\operatorname{{\rm svm}}(t)=\max(0,1-t)$. The SVM loss, also referred to as Hinge Loss, serves as a loss function in training classifiers, such as in Support Vector Machines. Following a similar approach to the one outlined above, we initiate the process by identifying a function $\beta(\cdot)$ that satisfies the properties stated in Lemma 20 for $\phi=\operatorname{{\rm svm}}$.
Lemma 34

### v2

@@ D.2 Coreset for Logistic Function
Proof.
Simplifying notation, let $\phi(t)=\operatorname{{\rm logistic}}(t)$. For a given $\alpha>0$, if we set $t=\alpha z$ and $K=\alpha^{2}k$,
@@
Up to this point, we have established $\sup_{t\in[0,\ \infty)}f_{K}(t)\leq 25+\frac{75K}{4\log K}$ under the condition that $K\geq e$. To conclude the argument, we need to determine an upper bound for $\sup_{t\in[0,\ \infty)}f_{K}(t)$ when $0<K\leq e$. Utilizing Lemma D.1, for $0<K\leq e$, we deduce
@@
Lemma D.9.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$. Then $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}_{\operatorname{{\rm logistic}}})$ and $(\mathcal{X},P,\bar{\mathcal{L}}_{\operatorname{{\rm logistic}}})$.
Proof.
As $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, we only need to determine $s(y)$ for $y$ with $\|y\|_{H}\leq A$. Henceforth, we assume that $\|y\|_{H}\leq A$. For $w\in\mathcal{W}=\{w\in\mathcal{X}\colon\|w\|_{H}\leq\frac{1}{A}\sqrt{\max(1,\log(A^{2}k))}\}$, Lemma D.8 indicates that Lemma D.2 is applicable with $B_{1}=A,B_{2}=\frac{1}{A}\sqrt{\max(1,\log(A^{2}k))},\phi(-B_{1}B_{2})\leq 1+\sqrt{\max(1,\log(A^{2}k))}=M$, and
@@
Using Lemma D.3 with $\gamma(\|x\|_{H})=\frac{\operatorname{{\rm logistic}}(0)}{M\beta_{\operatorname{{\rm logistic}}}(\|x\|_{H})}\geq\frac{1}{2M\beta_{\operatorname{{\rm logistic}}}(\|x\|_{H})}$, we conclude that
@@
So, $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ is an upper sensitivity function for $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$. The proof for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$ proceeds in a similar manner. ∎
Theorem D.10 (No Dimensional Logistic Coreset).
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is probability measure over $\mathcal{X}$ such that $P\left(\left\{x\in\mathcal{X}\colon\|x\|_{H}\geq A\right\}\right)=0$, $S=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$, and $C=(2A+1)\max(4Ak,2.5)+8Ak+1$. For $m\geq\frac{2S}{\varepsilon^{2}}(8C^{2}+S\log\frac{2}{\delta})$, any iid sample $x_{1},\ldots,x_{m}\in\mathcal{X}$ according to $P$ with weights $u_{i}=\frac{1}{m}$ provides an $\varepsilon$-coreset for $(\mathcal{X},P,\mathcal{X},\ell^{H}_{\operatorname{{\rm logistic}},k})$ (respectively for $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm logistic}},k})$ ) with probability at least $1-\delta$.
Proof.
Lemma D.9 asserts that $s(y)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ serves as an upper sensitivity function for $(\mathcal{X},\ P,\ \mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$ and $(\mathcal{X},P,\mathcal{X},\bar{\ell}^{H}_{\operatorname{{\rm logistic}},k})$. Since it is a constant function, $s$-sensitivity sampling is equivalent to sampling according to $P$. It is worth noting that $\operatorname{{\rm logistic}}$ is a $1$-Lipschitz, convex, and decreasing function. For $E_{1}=A$, we have
@@
This implies that Definition A.1 is satisfied for $\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k}$ and thus Theorem A.3 concludes the statement for $m\geq\frac{2S}{\varepsilon^{2}}\left(8C^{2}+S\log\frac{2}{\delta}\right)$. ∎
@@
Proof.
Because of similarity, we only handle the case $(\mathbb{R}^{d},P,\mathbb{R}^{d},\ell_{\operatorname{{\rm logistic}},k})$. Given that Lemma D.9 asserts the constant function $s(x)=1+\frac{340(1+kA^{2})}{\sqrt{\max(1,\log(kA^{2}))}}$ as an upper sensitivity function for $\mathcal{L}_{\operatorname{{\rm logistic}},k}$, we deduce $\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ)=\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$. To clarify, every function in $\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}}$ is a function in $\mathcal{L}_{\operatorname{{\rm logistic}},k}$ scaled by a positive constant. This implies ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{\mathcal{L}_{\operatorname{{\rm logistic}},k}},\succ))={\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)).$ We would like to remind that
@@
which concludes that $\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ)$ only includes half-spaces, empty set, and the whole space $\mathbb{R}^{d}$. Therefore, by Radon’s theorem, ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{L}_{\operatorname{{\rm logistic}},k},\succ))\leq d+1$ (for a proof, see Lemma 10.3.1 in  Matousek 2002). Leveraging Theorem 1.2, we derive the statement for $m=O(\frac{S}{\varepsilon^{2}}\left(d\log S+\log\frac{1}{\delta}\right))$, thereby completing the proof. ∎
In the previous theorem, we assumed that the data is hard-bounded. In the following results, we try to relax this assumption to a kind of soft bounding similar to what we have in Theorem D.6.
Lemma D.12.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$ and $P$ is a probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}^{2}.$ Then
@@
Proof.
As $\operatorname*{\mathbb{E}}_{x\sim P}(\|x\|_{H}^{2})\leq E^{2}_{1}$, Markov’s inequality implies $\operatorname*{\mathbb{P}}_{x\sim P}(\|x\|_{H}^{2}\geq 2E^{2}_{1})\leq\frac{1}{2}$. For an arbitrary $y\in\mathcal{X}$, define $\mathcal{X}_{y}=\left\{x\in\mathcal{X}\colon\|x\|_{H}\leq\sqrt{2}E_{1}\max(1,\frac{\|y\|_{H}}{E_{1}})\right\}$. Additionally, define
@@
Lemma D.8 indicates that Lemma D.2 is applicable with
@@
(Lemma D.2) $\leq\sup_{\|w\|_{H}\leq B_{2}}\frac{1}{\int_{x\in\mathcal{X}_{y}}\frac{\operatorname{{\rm logistic}}(0)}{M_{y}\beta(\|x\|_{H})}{\rm d}P(x)}$
@@
((*$\quad$( ∗ ) $\leq 2\left(2+\frac{\|y\|_{H}}{E_{1}}\right)\sqrt{\max(1,\log(2E_{1}^{2}k)}\frac{340(1+kE_{1}^{2})}{\max(1,\log(2kE_{1}^{2}))}.$
@@
To see (*), note that $M_{y}\leq\left(2+\frac{\|y\|_{H}}{E_{1}}\right)\sqrt{\max(1,\log(2E_{1}^{2}k)}$, $\operatorname{{\rm logistic}}(2)\geq 0.5$, and
@@
$\sup_{\|w\|_{H}\geq B_{2}}\frac{\ell^{H}_{{}_{\operatorname{{\rm logistic}},k}}(y,w)}{\int_{x\in\mathcal{X}}\ell^{H}_{{}_{\operatorname{{\rm logistic}},k}}(x,w){\rm d}P(x)}$ $\leq\sup_{\|w\|_{H}\geq B_{2}}\frac{\frac{1}{k}\|w\|_{H}^{2}+\log(1+e^{-\mathsf{K}(y,w)})}{\frac{1}{k}\|w\|_{H}^{2}}$
$\leq\sup_{\|w\|_{H}\geq B_{2}}\left[1+\frac{k}{\|w\|_{H}^{2}}\log(1+e^{-\mathsf{K}(y,w)})\right]$
@@
A similar argument concludes the statement for $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$. ∎
Theorem D.13.
Assume that $H$ is a reproducing kernel Hilbert space of real-valued functions on $\mathcal{X}$ with kernel $\mathsf{K}\colon\mathcal{X}\times\mathcal{X}\longrightarrow\mathbb{R}$, $P$ is probability measure over $\mathcal{X}$ such that $\operatorname*{\mathbb{E}}\limits_{x\sim p}\|x\|_{H}^{2}\leq E_{1}$,
@@
Proof.
Lemma D.12 indicates that $s(\cdot)$ and $S$ are an upper sensitivity function its corresponding total sensitivity for both $(\mathcal{X},P,\mathcal{L}^{H}_{\operatorname{{\rm logistic}},k})$ and $(\mathcal{X},P,\bar{\mathcal{L}}^{H}_{\operatorname{{\rm logistic}},k})$. Note that $\operatorname{{\rm logistic}}(\cdot)$ is decreasing and convex. using Equation (6), we have
@@
Now, the proof follows by applying Theorem C.6. ∎
Returning to the scenario where our Hilbert space is $\mathbb{R}^{d}$ equipped with the standard inner product as its kernel, to utilize Theorem 1.2, the computation of ${\rm VCdim}(\operatorname{\bf Ranges}(\mathcal{T}_{{\mathcal{L}_{\operatorname{{\rm logistic}},k}}},\succ))$ is necessary. The following theorem serves as a valuable instrument for handling the VC-dimension of linked-range spaces.
@@
Theorem D.14 (Anthony & Bartlett 2009, Theorem 8.14).
@@
The inequality $f_{w}(x)>r$ is not simply computable since it contains the evaluation $\|x\|$ (to compute $s(x)$), which cannot be done by the allowed operations. To circumvent this issue, we propose a modification to the upper sensitivity function $s$ outlined in Lemma D.12 as follows:
@@
So, using Theorems 1.2 and D.14, we have the next statement.
Theorem D.15.
@@
One might question the necessity of having Theorems D.10 and D.7 while Theorems D.13 and D.15 appear stronger and broader. It is crucial to note that in Theorems D.13 and D.15, the process demands re-sampling using $s$-sensitivity samples, which relies on evaluating the function $s$. In contrast, Theorems D.10 and D.7 permit the direct (via uniform sampling) use of given data points since they were originally sampled from $P$.
D.3 Coreset for svm Function
We would like to remind the reader that $\operatorname{{\rm svm}}(t)=\max(0,1-t)$. The SVM loss, also referred to as Hinge Loss, serves as a loss function in training classifiers, such as in Support Vector Machines. Following a similar approach to the one outlined above, we initiate the process by identifying a function $\beta(\cdot)$ that satisfies the properties stated in Lemma D.2 for $\phi=\operatorname{{\rm svm}}$.
Lemma D.16.